Francisco opened a saving bond that was compounded semiannually at 2.46%. If
he deposited $1,800, how much will be in the account after 5 years.

Found 2 solutions by Alan3354, mananth:
Answer by Alan3354(69443)

Francisco opened a saving bond that was compounded semiannually at 2.46%. If
he deposited $1,800, how much will be in the account after 5 years.
-------------
Assuming it's 2.46% per year:
5 years is 10 compounding periods.
FV = PV*(1 + 0.0246/2)^10 = 1800*(1.0123)^10
Just calculator work.

Answer by mananth(16946)


A = P + I where
P (principal) = $1,800.00
2.46% = 0246/2= 0.0123
5 years
2 times/year
A = P(1 + r/n)^nt
A = 1,800.00(1 + 0.0246/2)^(2)(5)
A = 1,800.00(1 + 0.0123)^(10)
A = $2,034.07